### v6

@@ 5 A cyclic group representation of relation algebra $34_{65}$
As an application, we give the first known finite representation of $34_{65}$. Relation algebra $34_{65}$ has four symmetric atoms $1^{\prime}$, $a$, $b$, and $c$, with forbidden cycles $bbc$ and $ccb$. The atom $a$ is flexible, hence $34_{65}$ is representable over a countable set.
@@ 6.1 $\mathcal{A}_{n}([i,i+j,i+\ell])$
If instead $j=\ell$, we only forbid $(a_{i+j},a_{i+j})$. Thus, $uv$ has $n^{2}-1$ needs in this case. Together with $u,v$, this yields $n^{2}+1$ points in total. ∎
Remark 6.2.
@@ 6.2 $33_{65}$
Consider the red edge $x_{2}x_{5}$, which has $9$ needs. The g-b need is met by $x_{0}$, and the g-g need is met by $x_{1}$. Now the r-r need is met by $x_{3},x_{4},x_{7},x_{8},x_{10}$. We can meet the r-g need with $x_{6}$ and an arbitrary need with $x_{9}$. This leaves $4$ unsatisfied needs, necessitating $4$ additional points. So we require at least $15$ points. ∎
@@
We have verified the unsatisfiability of each such $\Phi^{(n)}$ with a SAT solver. ∎
Remark 6.5.
@@ 6.3 $34_{65}$
  • $x_{2}$: $x_{2}x_{4}$, $x_{2}x_{6}$, $x_{2}x_{7}$, $x_{2}x_{8}$, $x_{2}x_{9}$ ($5$ edges)
  • $x_{3}:$ $x_{3}x_{6},x_{3}x_{7},x_{3}x_{8}$ ($3$ edges)
  • $x_{4}$: $x_{4}x_{5},x_{4}x_{6}$, $x_{4}x_{7},x_{4}x_{8}$ ($4$ edges)
  • $x_{5}$: $x_{5}x_{7},x_{5}x_{9}$ ($2$ edges)
  • $x_{6}$: $x_{6}x_{7},x_{6}x_{9}$ ($2$ edges)
@@
So there are two unmet needs, which necessitate two additional points. This brings our total to a minimum of $13$ points. ∎
@@
We have verified the unsatisfiability of each such $\Phi^{(n)}$ with a SAT solver. ∎
Remark 6.8.
@@ 7 Lower Bounds on Cyclic Group Representations of $33_{65}$
As $33_{65}$ is a symmetric relation algebra, we must enforce that the group representation is symmetric. Precisely, we must enforce that $i$ and $-i$ receive the same color (we are considering the additive group $\mathbb{Z}/n\mathbb{Z}$, and so $-i$ is the additive inverse of $i$). We enforce symmetry with the following constraint:
@@
Finally, we will define a Boolean formula to enforce that there are no forbidden cycles. Recall for $33_{65}$ that the forbidden cycles are of the form $ccc$ (green, green,green), $bcc$ (blue, green, green), and $cbb$ (green, blue, blue). To accomplish this, we explicitly enumerate the permissible cycles: any cycle containing the flexible atom $a$ (red), and the all-blue triangle $bbb$. Define:
@@ 8 Conclusion
Conjecture 8.1 (Strong Commutative Flexible Atom Conjecture).
@@
Remark 8.2.
The assumption that the relation algebras are commutative is indeed necessary, and we discovered this only after the conference version [AAL23] where we omitted the commutative hypothesis. The algebra $1284_{1316}$ has diversity atoms $a$, $b$, $r$, and $\breve{r}$, where $a$ is flexible and the only forbidden diversity cycle is $rbr$. However, $1284_{1316}$ is not commutative, and therefore not representable over an Abelian group. In particular, Comer’s method does not yield a group representation.
@@ Appendix A Point Algebra is not Finitely Representable
To see that no representation over a finite set is possible, we observe first that the interpretation of $r$ has to be a strict total ordering on $U$. Let $(u,v)\in r$. (We say $u$ is less than $v$.) Because $r$ is contained in $\breve{r}\circ r$, there is some point $w$ less than both $u$ and $v$. Similarly, because $r$ is contained in $r\circ\breve{r}$, there is some point $x$ greater than both $u$ and $v$. Hence the ordering has no endpoints, so cannot be finite. ∎

### v7

@@ 5 A cyclic group representation of relation algebra $34_{65}$
As an application, we give the first known finite representation of $34_{65}$. Relation algebra $34_{65}$ has four symmetric atoms $1^{\prime}$, $a$, $b$, and $c$, with forbidden cycles bbcbbc and ccbccb. The atom $a$ is flexible, hence $34_{65}$ is representable over a countable set.
@@ 6.1 $\mathcal{A}_{n}([i,i+j,i+\ell])$
If instead $j=\ell$, we only forbid $(a_{i+j},a_{i+j})$. Thus, $uv$ has $n^{2}-1$ needs in this case. Together with $u,v$, this yields $n^{2}+1$ points in total.
Remark 6.3
@@ 6.2 $33_{65}$
Consider the red edge $x_{2}x_{5}$, which has $9$ needs. The g-b need is met by $x_{0}$, and the g-g need is met by $x_{1}$. Now the r-r need is met by $x_{3},x_{4},x_{7},x_{8},x_{10}$. We can meet the r-g need with $x_{6}$ and an arbitrary need with $x_{9}$. This leaves $4$ unsatisfied needs, necessitating $4$ additional points. So we require at least $15$ points.
@@
We have verified the unsatisfiability of each such $\Phi^{(n)}$ with a SAT solver.
Remark 6.8
@@ 6.3 $34_{65}$
  • $x_{2}$: $x_{2}x_{4}$, $x_{2}x_{6}$, $x_{2}x_{7}$, $x_{2}x_{8}$, $x_{2}x_{9}$ (55 edges)
  • $x_{3}:$ $x_{3}x_{6},x_{3}x_{7},x_{3}x_{8}$ (33 edges)
  • $x_{4}$: $x_{4}x_{5},x_{4}x_{6}$, $x_{4}x_{7},x_{4}x_{8}$ (44 edges)
  • $x_{5}$: $x_{5}x_{7},x_{5}x_{9}$ (22 edges)
  • $x_{6}$: $x_{6}x_{7},x_{6}x_{9}$ (22 edges)
@@
So there are two unmet needs, which necessitate two additional points. This brings our total to a minimum of $13$ points.
@@
We have verified the unsatisfiability of each such $\Phi^{(n)}$ with a SAT solver.
Remark 6.13
@@ 7 Lower Bounds on Cyclic Group Representations of $33_{65}$
As $33_{65}$ is a symmetric relation algebra, we must enforce that the group representation is symmetric. Precisely, we must enforce that $i$ and i-i receive the same color (we are considering the additive group $\mathbb{Z}/n\mathbb{Z}$, and so i-i is the additive inverse of $i$). We enforce symmetry with the following constraint:
@@
Finally, we will define a Boolean formula to enforce that there are no forbidden cycles. Recall for $33_{65}$ that the forbidden cycles are of the form cccccc (green, green,green), bccbcc (blue, green, green), and cbbcbb (green, blue, blue). To accomplish this, we explicitly enumerate the permissible cycles: any cycle containing the flexible atom $a$ (red), and the all-blue triangle bbbbbb. Define:
@@ 8 Conclusion
Conjecture 8.1 (Strong Commutative Flexible Atom Conjecture)
@@
Remark 8.2
The assumption that the relation algebras are commutative is indeed necessary, and we discovered this only after the conference version [AAL23] where we omitted the commutative hypothesis. The algebra $1284_{1316}$ has diversity atoms $a$, $b$, $r$, and $\breve{r}$, where $a$ is flexible and the only forbidden diversity cycle is rbrrbr. However, $1284_{1316}$ is not commutative, and therefore not representable over an Abelian group. In particular, Comer’s method does not yield a group representation.
@@ Appendix A Point Algebra is not Finitely Representable
To see that no representation over a finite set is possible, we observe first that the interpretation of $r$ has to be a strict total ordering on $U$. Let $(u,v)\in r$. (We say $u$ is less than $v$.) Because $r$ is contained in $\breve{r}\circ r$, there is some point $w$ less than both $u$ and $v$. Similarly, because $r$ is contained in $r\circ\breve{r}$, there is some point $x$ greater than both $u$ and $v$. Hence the ordering has no endpoints, so cannot be finite.